# Fine-R1: Make Multi-modal LLMs Excel in Fine-Grained Visual Recognition by Chain-of-Thought Reasoning

### v1

@@ 1 Introduction
The visual world exhibits inherently fine-grained characteristics, which pose significant challenges for visual understanding. Objects are organized hierarchically according to shared traits and span a vast number of fine-grained categories. For example, the coarse-grained super-category bird can be further subdivided into thousands of fine-grained sub-categories such as Acadian Flycatcher, Great Crested Flycatcher, and Least Flycatcher. Moreover, new categories can emerge in real-world applications, requiring to identify unseen concepts (Geng et al., 2020). According to the latest statistics from the International Ornithologists’ Union (IOC), as of 2024, 11,276 bird species have been identified worldwide, and the number continues to grow with new species discovered.
Recent advances in Multi-modal Large Language Models (MLLMs) have achieved impressive results on general vision-language tasks such as image captioning and visual question answering (Liu et al., 2024a; b). However, prior studies (Zhang et al., 2024d; Liu et al., 2024c; He et al., 2025) reveal a substantial drop in performance when MLLMs are applied to knowledge-intensive fine-grained visual recognition (FGVR) task. FGVR, a long-standing challenge in computer vision, requires distinguishing subtle differences among visually similar categories, such as animal species (Wah et al., 2011), plant varieties (Nilsback and Zisserman, 2008), or specific models of cars (Krause et al., 2013) and aircraft (Maji et al., 2013). These tasks are difficult even for humans, as they demand extensive domain knowledge to resolve minimal inter-class variance and large intra-class variations. Notably, even state-of-the-art generative MLLMs such as GPT-4 (Achiam et al., 2023) and GeminiPro (Team et al., 2023) underperform compared to contrastive CLIP models (Radford et al., 2021; Zhai et al., 2023) dedicated for discriminative tasks.
To improve FGVR capability, early studies have explored fine-tuning MLLMs with classification data (Zhang et al., 2024d; He et al., 2025; Shi et al., 2025b). While this can yield gains, transforming general-purpose MLLMs into fine-grained classifiers requires extensive labeled data, which is costly to obtain. Moreover, these models tend to overfit to seen categories during training, limiting their utility in real-world scenarios where recognition of novel concepts is crucial (Geng et al., 2020).
@@ 2 Related Work
MLLMs for FGVR. Recent efforts to adapt MLLMs for classification either fine-tune them with classification data or design training-free approaches. Fine-tuning studies show that explicit object mentions in training data are crucial (Geigle et al., 2024; Zhang et al., 2024d), and that integrating sufficient classification-focused data enables MLLMs to bridge the gap between state-of-the-art classifiers while enhancing object-centric reasoning. Some works attribute underperformance to misalignment between visual objects and category names, addressing it through attribute-based alignment (He et al., 2025) or interpretable data synthesis (Shi et al., 2025a). In contrast, training-free methods such as Sparse Attention Vectors exploit sparse attention activations for discriminative tasks (Mitra et al., 2024a), but performance remains limited without large annotated datasets.
Reinforcement Learning. Reinforcement learning (RL) has shown strong potential in enhancing reasoning abilities of both LLMs and MLLMs by reducing reliance on large annotated datasets (Yue et al., 2024). In LLMs, the GPT series (Hurst et al., 2024; Achiam et al., 2023; Jaech et al., 2024) leveraged RL with human feedback (Ouyang et al., 2022) to optimize reasoning, excelling in mathematics (Cai et al., 2024; Ying et al., 2024; Shao et al., 2024; Yang et al., 2024; Luong et al., 2024) and coding tasks (Hui et al., 2024; Zhang et al., 2024a; e). DeepSeek-R1-Zero (Guo et al., 2025) further demonstrated RL-only training for reasoning improvement. In MLLMs, a series of work (Liu et al., 2025b; Li et al., 2025; Huang et al., 2025; Ma et al., 2025) advanced visual perception and inference through RL with verifiable rewards. Together, these works highlight RL’s effectiveness in interactive visual-linguistic reasoning. While policy optimization has shown great potential in image classification task (Liu et al., 2025b; Ma et al., 2025) for MLLMs, how it can be optimized for solving the key challenge of fine-grained image classification task hasn’t been investigated. Instead, we equip policy optimization with contrastive paradigms to make the model more robust to the high intra-class variance and discriminative to the low inter-class variance.
@@ 5.1 Experiment Settings
Evaluation Metrics. We define success on a single example as whether the ground-truth choice is included in the MLLM generation. We report the success rate of all test examples as the accuracy in the closed-world setting. Since evaluating models in the open-world setting presents additional challenges, as predictions may differ in granularity (e.g., Boeing 737 vs. Boeing 737-200), or ground truth may include redundancy for distinguishing from others (e.g.,“Coupe 2012” in Audi A5 Coupe 2012 and Audi S5 Coupe 2012), we use two complementary metrics: (1) text inclusion (Zhang et al., 2024d), evaluating strict string matching. (2) relative semantic similarity between the text embeddings of predictions and ground truth calculated by the SigLIP (Zhai et al., 2023) text encoder. Instead of using the similarity as reward directly, we use the similarity between the super-category and the ground truth subcategory as the standard to calculate the relative value. Formally, the relative semantic similarity $SS_{\text{relative}}$ is expressed as:
@@ 5.3 Ablation Study
Inference Strategies. We investigate two inference strategies, including CoT prompting, and In-Context Learning (ICL) in the closed-world evaluation. For CoT prompting, we leverage the zero-shot CoT prompting technique by adding “let’s think step by step” at the end of the prompt (Wei et al., 2022a; Kojima et al., 2022). For CLIP-like models, we additionally add prompt ensembling results for SigLIP using 80 prompt templates from ImageNet dataset. As shown in Figure 3(b), compared to the baseline (74.68%), direct CoT prompting without training for CoT reasoning (74.79%) has a limited impact on FGVR, which is also affirmed in (Zhang et al., 2024c). For ICL, we randomly sample one demonstration for each candidate once it belongs to seen categories (i.e., occurs in the training data). However, since we can only retrieve demonstrations for seen categories in the candidates, the context may introduce bias. The results show that Fine-R1-3B surpasses Qwen2.5-VL-3B with ICL by 17.90%, strengthen the effectiveness of Fine-R1. It is worth noting that Fine-R1 outperforms CLIP-like models even with prompt ensembling optimization.
@@ Appendix F General Capability
To comprehensively assess the model’s general capabilities endowed with FGVR capability, we conduct evaluations on two set of datasets: (1) classification-based VQA benchmark: ImageWikiQA (Zhang et al., 2024d), which is a multiple-choice question-answering dataset collected by feeding the Wikipedia pages of ImageNet classes to GPT-4. (2) General VQA benchmarks: MME (Fu et al., 2023a), MMBench (Liu et al., 2024d), and SEED-Bench (Li et al., 2023a). As shown in Table 7, we find that current MLLMs perform poorly in answering these questions, suggesting that their poor FGVR performance is a fundamental limitation for more advanced capabilities. However, Fine-R1 raises the performance from 54.85% to 58.45%, demonstrating that FGVR is indeed a foundation for MLLMs’ advanced capabilities. Moreover, Fine-R1 demonstrates competitive in general-purpose performance and even achieves improvements on MMBench and SEED-Bench. It is worth noting that Fine-R1 is post-trained solely on the FGVR task without incorporating general instruction tuning data, proving that RL drives performance gains not by mere answer memorization. These results suggest that Fine-R1 can serve both as a specialized assistant for users interested in FGVR and as a general-purpose MLLM for broader applications.

### v2

@@ 1 Introduction
The visual world exhibits inherently fine-grained characteristics, which pose significant challenges for visual understanding. Objects are organized hierarchically according to shared traits and span a vast number of fine-grained categories (Zhang et al., 2024c). For example, the coarse-grained super-category bird can be further subdivided into thousands of fine-grained sub-categories such as Acadian Flycatcher, Great Crested Flycatcher, and Least Flycatcher. Moreover, new categories can emerge in real-world applications, requiring to identify unseen concepts (Geng et al., 2020). According to the latest statistics from the International Ornithologists’ Union (IOC), as of 2024, 11,276 bird species have been identified worldwide, and the number continues to grow with new species discovered.
Recent advances in Multi-modal Large Language Models (MLLMs) have achieved impressive results on general vision-language tasks such as image captioning and visual question answering (Liu et al., 2024a; b). However, prior studies (Zhang et al., 2024e; Liu et al., 2024c; He et al., 2025) reveal a substantial drop in performance when MLLMs are applied to knowledge-intensive fine-grained visual recognition (FGVR) task. FGVR, a long-standing challenge in computer vision, requires distinguishing subtle differences among visually similar categories, such as animal species (Wah et al., 2011), plant varieties (Nilsback and Zisserman, 2008), or specific models of cars (Krause et al., 2013) and aircraft (Maji et al., 2013). These tasks are difficult even for humans, as they demand extensive domain knowledge to resolve minimal inter-class variance and large intra-class variations. Notably, even state-of-the-art generative MLLMs such as GPT-4 (Achiam et al., 2023) and GeminiPro (Team et al., 2023) underperform compared to contrastive CLIP models (Radford et al., 2021; Zhai et al., 2023) dedicated for discriminative tasks.
To improve FGVR capability, early studies have explored fine-tuning MLLMs with classification data (Zhang et al., 2024e; He et al., 2025; Shi et al., 2025b). While this can yield gains, transforming general-purpose MLLMs into fine-grained classifiers requires extensive labeled data, which is costly to obtain. Moreover, these models tend to overfit to seen categories during training, limiting their utility in real-world scenarios where recognition of novel concepts is crucial (Geng et al., 2020).
@@ 2 Related Work
MLLMs for FGVR. Recent efforts to adapt MLLMs for FGVR either fine-tune them with classification data or design training-free approaches (Peng et al., 2025). Fine-tuning studies show that explicit object mentions in training data are crucial (Geigle et al., 2024; Zhang et al., 2024e), and that integrating sufficient classification-focused data enables MLLMs to bridge the gap between state-of-the-art classifiers while enhancing object-centric reasoning. Some works attribute underperformance to misalignment between visual objects and category names, addressing it through attribute-based alignment (He et al., 2025) or interpretable data synthesis (Shi et al., 2025a). In contrast, training-free methods such as Sparse Attention Vectors exploit sparse attention activations for discriminative tasks (Mitra et al., 2024a), but performance remains limited without large annotated datasets.
Reinforcement Learning. Reinforcement learning (RL) has shown strong potential in enhancing reasoning abilities of both LLMs and MLLMs by reducing reliance on large annotated datasets (Yue et al., 2024). In LLMs, the GPT series (Hurst et al., 2024; Achiam et al., 2023; Jaech et al., 2024) leveraged RL with human feedback (Ouyang et al., 2022) to optimize reasoning, excelling in mathematics (Cai et al., 2024; Ying et al., 2024; Shao et al., 2024; Yang et al., 2024; Luong et al., 2024) and coding tasks (Hui et al., 2024; Zhang et al., 2024a; f). DeepSeek-R1-Zero (Guo et al., 2025) further demonstrated RL-only training for reasoning improvement. In MLLMs, a series of work (Liu et al., 2025b; Li et al., 2025; Huang et al., 2025; Ma et al., 2025) advanced visual perception and inference through RL with verifiable rewards. Together, these works highlight RL’s effectiveness in interactive visual-linguistic reasoning. While policy optimization has shown great potential in image classification task (Liu et al., 2025b; Ma et al., 2025) for MLLMs, how it can be optimized for solving the key challenge of fine-grained image classification task hasn’t been investigated. Instead, we equip policy optimization with contrastive paradigms to make the model more robust to the high intra-class variance and discriminative to the low inter-class variance.
@@ 5.1 Experiment Settings
Evaluation Metrics. We define success on a single example as whether the ground-truth choice is included in the MLLM generation. We report the success rate of all test examples as the accuracy in the closed-world setting. Since evaluating models in the open-world setting presents additional challenges, as predictions may differ in granularity (e.g., Boeing 737 vs. Boeing 737-200), or ground truth may include redundancy for distinguishing from others (e.g.,“Coupe 2012” in Audi A5 Coupe 2012 and Audi S5 Coupe 2012), we use two complementary metrics: (1) text inclusion (Zhang et al., 2024e), evaluating strict string matching. (2) relative semantic similarity between the text embeddings of predictions and ground truth calculated by the SigLIP (Zhai et al., 2023) text encoder. Instead of using the similarity as reward directly, we use the similarity between the super-category and the ground truth subcategory as the standard to calculate the relative value. Formally, the relative semantic similarity $SS_{\text{relative}}$ is expressed as:
@@ 5.3 Ablation Study
Inference Strategies. We investigate two inference strategies, including CoT prompting, and In-Context Learning (ICL) in the closed-world evaluation. For CoT prompting, we leverage the zero-shot CoT prompting technique by adding “let’s think step by step” at the end of the prompt (Wei et al., 2022a; Kojima et al., 2022). For CLIP-like models, we additionally add prompt ensembling results for SigLIP using 80 prompt templates from ImageNet dataset. As shown in Figure 3(b), compared to the baseline (74.68%), direct CoT prompting without training for CoT reasoning (74.79%) has a limited impact on FGVR, which is also affirmed in (Zhang et al., 2024d). For ICL, we randomly sample one demonstration for each candidate once it belongs to seen categories (i.e., occurs in the training data). However, since we can only retrieve demonstrations for seen categories in the candidates, the context may introduce bias. The results show that Fine-R1-3B surpasses Qwen2.5-VL-3B with ICL by 17.90%, strengthen the effectiveness of Fine-R1. It is worth noting that Fine-R1 outperforms CLIP-like models even with prompt ensembling optimization.
@@ Appendix F General Capability
To comprehensively assess the model’s general capabilities endowed with FGVR capability, we conduct evaluations on two set of datasets: (1) classification-based VQA benchmark: ImageWikiQA (Zhang et al., 2024e), which is a multiple-choice question-answering dataset collected by feeding the Wikipedia pages of ImageNet classes to GPT-4. (2) General VQA benchmarks: MME (Fu et al., 2023a), MMBench (Liu et al., 2024d), and SEED-Bench (Li et al., 2023a). As shown in Table 7, we find that current MLLMs perform poorly in answering these questions, suggesting that their poor FGVR performance is a fundamental limitation for more advanced capabilities. However, Fine-R1 raises the performance from 54.85% to 58.45%, demonstrating that FGVR is indeed a foundation for MLLMs’ advanced capabilities. Moreover, Fine-R1 demonstrates competitive in general-purpose performance and even achieves improvements on MMBench and SEED-Bench. It is worth noting that Fine-R1 is post-trained solely on the FGVR task without incorporating general instruction tuning data, proving that RL drives performance gains not by mere answer memorization. These results suggest that Fine-R1 can serve both as a specialized assistant for users interested in FGVR and as a general-purpose MLLM for broader applications.